Week of 170508

WLCG Operations Call details

  • At CERN the meeting room is 513 R-068.

  • For remote participation we use the Vidyo system. Instructions can be found here.

General Information

  • The purpose of the meeting is:
    • to report significant operational issues (i.e. issues which can or did degrade experiment or site operations) which are ongoing or were resolved after the previous meeting;
    • to announce or schedule interventions at Tier-1 sites;
    • to inform about recent or upcoming changes in the experiment activities or systems having a visible impact on sites;
    • to provide important news about the middleware;
    • to communicate any other information considered interesting for WLCG operations.
  • The meeting should run from 15:00 until 15:20, exceptionally to 15:30.
  • The SCOD rota for the next few weeks is at ScodRota
  • General information about the WLCG Service can be accessed from the Operations Web
  • Whenever a particular topic needs to be discussed at the daily meeting requiring information from site or experiments, it is highly recommended to announce it by email to wlcg-operations@cernSPAMNOTNOSPAMPLEASE.ch to make sure that the relevant parties have the time to collect the required information or invite the right people at the meeting.

Tier-1 downtimes

Experiments may experience problems if two or more of their Tier-1 sites are inaccessible at the same time. Therefore Tier-1 sites should do their best to avoid scheduling a downtime classified as "outage" in a time slot overlapping with an "outage" downtime already declared by another Tier-1 site supporting the same VO(s). The following procedure is recommended:

  1. A Tier-1 should check the downtimes calendar to see if another Tier-1 has already an "outage" downtime in the desired time slot.
  2. If there is a conflict, another time slot should be chosen.
  3. In case stronger constraints cannot allow to choose another time slot, the Tier-1 will point out the existence of the conflict to the SCOD mailing list and at the next WLCG operations call, to discuss it with the representatives of the experiments involved and the other Tier-1.

As an additional precaution, the SCOD will check the downtimes calendar for Tier-1 "outage" downtime conflicts at least once during his/her shift, for the current and the following two weeks; in case a conflict is found, it will be discussed at the next operations call, or offline if at least one relevant experiment or site contact is absent.

Links to Tier-1 downtimes




  • local: Andrei (databases), Jesus (storage), Julia (WLCG), Maarten (SCOD + ALICE + GGUS), Marian (networks), Paul (storage), Tomas (ATLAS), Vincent (security)
  • remote: Alberto (monitoring), Antonio (LHCb), David (FNAL), Di (TRIUMF), Dimitri (KIT), Dmytro (NDGF), John (RAL), Pepe (PIC), Sang-Un (KISTI), Tommaso (CMS)

Experiments round table:

  • ATLAS reports ( raw view) -
    • Activities: Reprocessing campaign is basically done. MC16 campaign runs in full speed. First version of the derivations also ran last week on the reprocessed data to provide recommendations to physics groups.
    • Issues (no big issues):
      • Deletion at StoRM sites, waiting on solution: (GGUS:126896)
      • Deletion at EOS (CERN-PROD) - space reporting issues. Fixed by EOS experts: the automatic remediation in place for a misbehaving bestman daemon.

  • CMS reports ( raw view) -
    • Magnet ramped, subdetectors tests + cosmic runs
    • activity on the low side
      • Another bunch of 2016 data rereco just submitted
      • for the rest, mostly production of 2017 trigger MC samples for HLT finalization
      • analysis steady at > 50 kcores
      • HLT back in the game after EOS authentication change
    • We experienced some problems in DBS APIs, still unclear whether we can wait for next monthly update or we need an unscheduled deployment of CMSWEB
    • KIBANA: since a few days CREAM and VOMS @ 50%
      • experts stated it is a fake alarm (GGUS:128146 and GGUS:128127), still it raises shifters attention
        • if indeed fake (and we have no reason to doubt it, since we see no real problem), they should be fixed
          • Marian: VOMS is OK now, CREAM is being looked into
    • few site level GGUS tickets, nothing which needs special attention.

  • ALICE -
    • Very high activity

  • LHCb reports ( raw view) -
    • Activity
      • Stripping v28 over, Stripping v24 waiting for developers.
    • Site Issues
      • T1:
        • RAL: Disk server failure last week, back in production today. Some FTS3 timeouts during staging.

Sites / Services round table:

  • BNL: AFS phased out from Tier1 WNs
  • CNAF:
    • We'll update table related to GGUS:128172 as soon as we finish IPV6 tests.
  • EGI:
  • IN2P3:
  • JINR:
  • KIT:
    • ARC CE instabilities are being investigated
  • NL-T1:
    • Unable to dial in due to agenda conflict
    • A pool node in the SURFsara dCache lost connection to its storage, GGUS:128189. Fixed by rebooting.
  • NRC-KI:
  • OSG:
  • PIC: NTR
  • RAL: We plan to upgrade Castor at RAL tomorrow (9th May) and Thursday (11th May). The central services and LHCb will be done tomorrow. Atlas and CMS will be done on Thursday.

  • CERN computing services:
  • CERN storage services: NTR
  • CERN databases: NTR
  • Monitoring:
    • Draft reports for the April availability distributed. Waiting for comments.
  • MW Officer: NTR
  • Networks: 181 perfSONARs updated to version 4.0 (out of 246)
  • Security: NTR


Edit | Attach | Watch | Print version | History: r15 < r14 < r13 < r12 < r11 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r15 - 2017-05-08 - MaartenLitmaath
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback